Tohban Report 2014-01-08

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 02 Jan 2014
End Date: 08 Jan 2014
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

Solar activity was relatively high this week, with one X-class flare and 11 M-class flares. AR 1944 produced many of these flares. It is still located near the disk center and is expected to produce more flares during the next few days. There are currently 8 ARs on the disk. STEREO B shows no significantly bright spots that would rotate onto the disk soon.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    39    11     1

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    15     8     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     6     2     0

There were RHESSI flares/GOES flares 231 / 51 over the time range '01-Jan-14 00:00', '08-Jan-14 16:00'

Memory Management

The SSR fill level was relatively high with maximum 41%. This high level is due to the relatively high solar activity and also partly due to an operator error omitting a BGS pass.

Spacecraft Status

The cold plate 1 and 2 are at about 118.6 K and 117 K, respectively. Cryocooler is nominal. Spin rate is decreasing and we may need a spin up in a couple of week.

Data Gaps


GAP START TIME              GAP END TIME                   GAP (SEC)
2014-01-02T22:10:00.000 -- 2014-01-03T00:00:00.000       6600.0000
2014-01-03T00:00:00.000 -- 2014-01-03T01:20:00.000       4800.0000

Detector issues

On Jan 3, Hugh Hudson noticed that the attenuator system has been in state 3. It turned out that, in Detector 6 (both front and rear), livetime was decreasing rapidly and preamp-reset and FAST LLD were also rapidly increasing. D6 was immediately removed from the attenuator calculation (It was replaced with D9 and the calculations are now being performed with detectors 1,3,7 & 9). Then, over the next few days, several adjustments were made to D6 high voltage and FAST threshold (both front and rear) as described in the Detector Update page. The monitor rates seem to have stabilized, as of this writing, and the next tohban will be analyzing background energy spectra to better understand the D6 behavior.

Other notes

There was an isolated spike in Detector 4 (monitor rates and summary count rates) at DOY 002 (Jan 2) 18:40 UT.

Spacecraft Management

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? None
Personal tools